Solve using the principle of zero products.
(21 - 0.8x)(5 - 0.4x) = 0

Answer by unlockmath(1688)

Hello,
Since this has been factored out (21 - 0.8x)(5 - 0.4x) = 0 we can solve both x's.
(21 - 0.8x)=0 Subtract 21 from both sides to give us:
-.8x=-21 Divide both sides by -.8 to be:
x=26.25 Now let's do the other x
(5 - 0.4x) = 0 Subtract 5 from both sides to be:
-.4x=-5 Divide both sides by -.4 to be:
x=12.5
Plug these answers into the original to see if it works.
RJ
Check out a new book I wrote at:
www.math-unlock.com